OpenStudy (anonymous):

write in standard form f(3,8) d: y=4

OpenStudy (anonymous):

ax+by=c general form. We start in y=mx+b form and so we get such an answer. 3=4(8)+b 3=32+b -29=b y=4x-29 -4x+y=-29 4x-y=29

OpenStudy (anonymous):

hope it'll help

OpenStudy (anonymous):

so wat is it written in standard form? im lost

OpenStudy (anonymous):

yes y=mx+b is the standard form

OpenStudy (anonymous):

i kno tht but i dnt know where to plug my numbers in at

OpenStudy (anonymous):

4x-y=29 <= this is the final answer i solved every thing step by step

OpenStudy (anonymous):

hmmm ohk